OpenStudy (anonymous):

How do you factor a binomial with a odd exponent. This format (X^3)-8

OpenStudy (unklerhaukus):

can you write 8 as a cube ?

OpenStudy (anonymous):

it's a find the limit problem thee numerator is (x-2) the denominator is the format

OpenStudy (unklerhaukus):

if you get (x^3)-8 into the form of a difference of cubes you will be able to factorise , one of the factors wil be (x-2)

OpenStudy (anonymous):

can you play this otu for me?

OpenStudy (unklerhaukus):

Sum of cubes\[a^3 + b^3 = (a + b)(a^2 – ab + b^2) \]Difference of cubes\[a^3 – b^3 = (a – b)(a^2 + ab + b^2)\]

OpenStudy (anonymous):

ah okay thx i am going to try that

OpenStudy (unklerhaukus):

hows it going?